# Transferable text data distillation by trajectory matching

### v1

@@ 5 Analysis
5.1 Influence of selected data
@@
RAND 5% 52.5
NACD 5% 53.7
To demonstrate the generalizability of our data distillation method and examine the impact of selected data, we replaced LESS with random sampling for data selection in this section. The OPT 1.3B experimental results when full dataset is $All^{*}$ are shown in Table 3. For randomly selected samples, NACD also brings improvements to ARC-Easy experimental results, which demonstrating our approach has few dependencies on specific data selection methods and opening up a new perspective for data compression LLM training.
5.2 Influence of the regularization loss
@@ 5.3 The computation cost
Table 5 describes the computation complexity and the actual time of NACD. Taking the OPT-1.3B experiment of ARC-Easy in Flan-v2 as an example, we can see the main computation time costs in expert trajectory and data distillation steps. Noteworthy is the fact that in practical implementation, once synthesized data is generated through the initial two steps, it becomes reusable for various models (section 4.2). This one-time effort in data synthesis proves to be a cost-effective approach, as the data distillation paradigm significantly reduces data volume, leading to substantial savings in subsequent instruction
@@ 5.4 Visualization of the synthesized data
The synthetic data obtained from NACD distillation is pseudo embedding. In order to observe the semantics of the synthesized data, we take the distillation experiment with Flan v2 on ARC-Easy as an example, select two samples for analysis and decode the nearest pseudo-tokens as shown in the Figure 3. It can be observed that the synthesized prompt data tends to provide hints towards the correct answers, such as "disaster" and "suspected" for negative reviews recognition.

### v2

@@ 5 Analysis
5.1 Influence of the selected data and the statistical significance
@@
RAND 5% 52.7(0.2)
NACD 5% 53.6(0.4)
To demonstrate the generalizability of our data distillation method and examine the impact of selected data, we replaced LESS with random sampling for data selection in this section. Besides this, to prove the statistical significance of our method, we have added experimental results of randomly selected data. Following the setting in [9], we ran the experiment three times with different seeds and provided the error bars.The OPT 1.3B experimental results when full dataset is $All^{*}$ are shown in Table 3. For randomly selected samples, NACD also brings improvements to ARC-Easy experimental results, which demonstrating our approach has few dependencies on specific data selection methods and opening up a new perspective for data compression LLM training.
5.2 Ablation studies
@@
The compression ratio was a important factor in NACD, in order to verify the stability of NACD in different scenarios, under the setting of "ARC-Easy+$All^{*}$", we added experiments with larger compression rates of 10% and smaller compression rates of 0.02%. As shown in Table 5, at different compression rates, NACD consistently shows improvement, further confirming the effectiveness of our method. Additionlly, we have added an ablation study on prompt length k and regularization coefficient $\beta$ in Table 6 . From the experimental results, it can be observed that, firstly, for different values of beta (100, 50, and 20), NACD shows stable performance. While we also attempted to increase k to 5, and even NACD exhibited better performance. However, when we increased k to 10, NACD started showing disastrous performance. We believe that the increase of the learning space has led to an excessive level of difficulty in learning. Therefore, in NACD, there exists a balance between the search space of data and the learning difficulty.
Table 5: Different compression ratios experimental results of the instruction tuning tasks in ARC-Easy. Bold numbers denote that our proposed method outperforms the baselines. NACD shows stable improvement in LESS under different compression ratios.
Base Model Method Full dataset Data % Accuracy
OPT-1.3B baseline $All^{*}$ 0% 51.1
FULL $All^{*}$ 100% 54.0
LESS $All^{*}$ 10% 53.2
LESS+NACD $All^{*}$ 10% 53.8
LESS $All^{*}$ 0.02% 51.1
LESS+NACD $All^{*}$ 0.02% 51.6
Table 6: Ablation studies of the instruction tuning tasks in ARC-Easy. $k$ denotes the prompt token length and $\beta$ denotes the weight factor in Equation 3. To specifically validate the effectiveness of the NACD method on other source data, we replaced the target full dataset with Dolly.
Base Model Method Full dataset $\beta$ $k$ Data % Accuracy
OPT-1.3B baseline Dolly - - 0% 51.1
FULL Dolly - - 100% 54.0
LESS Dolly - - 5% 50.8
LESS+NACD Dolly 100 5 5% 52.0
LESS+NACD Dolly 100 2 5% 51.0
LESS+NACD Dolly 50 2 5% 51.2
LESS+NACD Dolly 20 2 5% 51.3
@@ 5.3 The computation cost
Table 7 describes the computation complexity and the actual time of NACD. Taking the OPT-1.3B experiment of ARC-Easy in Flan-v2 as an example, we can see the main computation time costs in expert trajectory and data distillation steps. Noteworthy is the fact that in practical implementation, once synthesized data is generated through the initial two steps, it is important to note that the distilled dataset obtained at once can be used multiple times for model training, and even for instruction learning tasks in different language models (the embedding layer inconsistent model in Section 4.2). This one-time effort in data synthesis proves to be a cost-effective approach, as the data distillation paradigm significantly reduces data volume, leading to substantial savings in subsequent instruction
@@ 5.4 Visualization of the synthesized data
The synthetic data obtained from NACD distillation is pseudo embedding. In order to observe the semantics of the synthesized data, we take the distillation experiment with Flan v2 on ARC-Easy as an example, select two samples for analysis and decode the nearest pseudo-tokens as shown in the Figure 3. It can be observed that the synthesized prompt data tends to provide hints towards the correct answers, such as "disaster" and "suspected" for negative reviews recognition. Additionlly